Still having some problems with my homework. Can anyone help?
9x-2y=50
6x-2y=32
Thank you,
Amanda
: Still having some problems with my homework. Can anyone help?
9x-2y=50
6x-2y=32
Thank you,
Amanda

Answer by jim_thompson5910(9401)


Start with the given system of equations:
system(9x-2y=50,6x-2y=32)


-1(6x-2y)=-1(32) Multiply the both sides of the second equation by -1.


-6x+2y=-32 Distribute and multiply.


So we have the new system of equations:
system(9x-2y=50,-6x+2y=-32)


Now add the equations together. You can do this by simply adding the two left sides and the two right sides separately like this:


(9x-2y)+(-6x+2y)=(50)+(-32)


(9x+-6x)+(-2y+2y)=50+-32 Group like terms.


3x+0y=18 Combine like terms. Notice how the y terms cancel out.


3x=18 Simplify.


x=(18)/(3) Divide both sides by 3 to isolate x.


x=6 Reduce.


------------------------------------------------------------------


9x-2y=50 Now go back to the first equation.


9(6)-2y=50 Plug in x=6.


54-2y=50 Multiply.


-2y=50-54 Subtract 54 from both sides.


-2y=-4 Combine like terms on the right side.


y=(-4)/(-2) Divide both sides by -2 to isolate y.


y=2 Reduce.


So our answer is x=6 and y=2.


Which form the ordered pair .


This means that the system is consistent and independent.


Notice when we graph the equations, we see that they intersect at . So this visually verifies our answer.


drawing(500,500,-4,16,-8,12,<BR>
grid(1),<BR>
graph(500,500,-4,16,-8,12,(50-9x)/(-2),(32-6x)/(-2)),<BR>
circle(6,2,0.05),<BR>
circle(6,2,0.08),<BR>
circle(6,2,0.10)<BR>
) Graph of 9x-2y=50 (red) and 6x-2y=32 (green)
Question 153878: Still having some problems with my homework. Can anyone help?
9x-2y=50
6x-2y=32
Thank you,
Amanda
: Still having some problems with my homework. Can anyone help?
9x-2y=50
6x-2y=32
Thank you,
Amanda

Answer by josmiceli(2035)

You have 2 equations and 2 unknowns, so it's solvable
9x - 2y = 50
6x - 2y = 32
Multiply both sides of the 2nd equation by -1
9x - 2y = 50
-6x + 2y = -32
Now add the equations
3x = 18
x = 6 answer
Substitute this value in the 1st equation
9x - 2y = 50
9*6 - 2y = 50
54 - 2y = 50
2y = 54 - 50
2y = 4
y = 2 answer
------------------
check answers:
9x - 2y = 50
6x - 2y = 32
-----------------
9*6 - 2*2 = 50
54 - 4 = 50
50 = 50
OK
6*6 - 2*2 = 32
36 - 4 = 32
32 = 32
OK